OpenStudy (anonymous):

A pole-vaulting student (mass 62 kg) clears a bar and then falls from rest from a height of 4.50 m onto a foam rubber pad below. What is her momentum at the moment of impact with the pad below?

OpenStudy (anonymous):

h=gtt/2 free fall u=0 t1=sq rt 2h/g v=u+at v=gt1 here momentum=mv